Three student teams collected 65 books for the school library. The first team collected 10 fewer books than the second team and the third team collected 30% of the number of books collected by first and second teams together. How many books did each team collect?

Answer by jorel1380(3719)

Let n be the number of books collected by the second team. Then the first team would be n-10; and the third team would be .3(n+n-10). So:
n+n-10+.3(2n-10)=65
2.6n=78
n=30
n-10=20
.3(2n-10)=15
The first team collected 20 books; the second team collected 30; and the third team collected 15. ☺☺☺☺